For what value of b will the line y=-2x+b be tangent to the parabola y=3x^2+4x-1
Answer by ikleyn(52781)

.
For what value of b will the line y=-2x+b be tangent to the parabola y=3x^2+4x-1
~~~~~~~~~~~~~~


            There are  TWO  WAYS  to solve the problem:  one way is  Algebra,  the other way is  Calculus.

            I will show you both.


1.   Calculus way


The slope of the given line is -2, the constant value.


The slope of the parabola is  6x+4  at the point with abscissa x.


In order for the line be tangent to the parabola, the necessary condition is "slope" = "slope",

which gives  -2 = 6x + 4,  and then  x= -1.


At x= -1,  the y-coordinate of the parabola is  3*(-1)^2 + 4*(-1) - 1 = 3*1 - 4 - 1 = -2;

           the y-coordinate of the line     is -2*(-1) + b= 2 + b.


Y-coordinate should be the same, which gives  2 + b = -2, and hence  b= -4.    ANSWER

So,  the  Calculus solution is completed.


2. Algebra way


You write this equation

    -2x + b = 3x^2 + 4x-1


to find common point of the line and the parabola.


Its standard form is

    3x^2 + 6x - (1+b) = 0.


The straight line is a tangent to the parabola if and only if there is ONLY ONE common point;

in other words, if there is ONLY ONE solution to the last equation.


It means that the discriminant of this equation is zero

    6^2 + 4*3*(1+b) = 0,

or

    36 + 12*(1+b) = 0

     3 + (1+b) = 0

     b         = -4.


We got the same answer.

Thus the problem is solved and two basic solutions are presented.